OpenStudy (anonymous):

2/3+4/5x=1/2-x solve. please show work

OpenStudy (anonymous):

2/3+4/5x=1/2-x 2/3-1/2=-x +4/5x Find the common denominator to add/subtract

OpenStudy (anonymous):

answer should be {-5/54} but i dont know how

OpenStudy (anonymous):

Do you understand what to do from what i left of?

OpenStudy (anonymous):

i dont understand what u mean

OpenStudy (anonymous):

Did you understand what I did?

OpenStudy (anonymous):

could u please work it out for me